# Fourier-Mukai Partners of Abelian Varieties and K3 surfaces in Positive and Mixed Characteristics

### v1

@@ 5.1. Crystalline cohomology
If $X$ and $Y$ are derived-equivalent $K3$ surfaces over $k$, they have the same Newton polygons. In particular, ordinarity and supersingularity of K3 surfaces are derived-invariant.
@@
where the ranks of the filtered pieces increase as $0,1,21,22$ The perfect Poincaré pairing on $H^{2}_{crys}(X/W(k))$ respects the $F$-crystal structure, so we see $Fil_{\rm slope}^{1}=(Fil_{\rm slope}^{2})^{\perp}$. We also endow the (even-degree) Mukai $F$-isocrystal
@@ 5.2. De Rham cohomology
where the ranks of the filtered pieces increase as $0,1,21,22$ Using the perfect duality, we have $Fil^{1}_{\rm Hdg}=(Fil_{\rm Hdg}^{2})^{\perp}$. We also obtain from the above that
@@ 5.3. Proofs of the main results for K3 surfaces
[…]
@@
Since $E$ defines a derived-equivalence $\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(X)\simeq\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(Y)$, we have
@@
A weaker version of the above can be obtained as follows as well. Let $X,Y$ be derived-equivalent K3 surfaces over $k$. We know from [LO15] that $Y\simeq M_{v}(X)$ can be written as a certain fine moduli space of stable sheaves on $X$. Then, the relative moduli space $\mathcal{M}_{v}(X_{\operatorname{can}})$ over $W(k)$ is fine and Fourier-Mukai equivalent to $X_{\operatorname{can}}$ via the universal stable sheaf (Corollary 2.7). Proposition 5.5 implies that $Y_{\operatorname{can}}\simeq\mathcal{M}_{v}(X_{\operatorname{can}})$, so $X_{\operatorname{can}},Y_{\operatorname{can}}$ are relative Fourier-Mukai partners. This argument, however, does not show that arbitrary Fourier-Mukai kernels can be lifted.
@@ 6. Sample application to a result in [HLT20]
    $A$ is a degree 2 étale cyclic cover of $E\times_{k}F$.
@@
    $A$ is a degree 3 étale cyclic cover of $E\times_{k}F$, and $F$ has an automorphism of order 3.
@@ Appendix A Crystalline and de Rham Chern classes
[…]

### v2

@@ 5.1. Crystalline cohomology
If $X$ and $Y$ are derived equivalent $K3$ surfaces over $k$, they have the same Newton polygons. In particular, ordinarity and supersingularity of K3 surfaces are derived-invariant.
@@
where the ranks of the filtered pieces increase as $0,1,21,22.$ The perfect Poincaré pairing on $H^{2}_{crys}(X/W(k))$ respects the $F$-crystal structure, so we see $Fil_{\rm slope}^{1}=(Fil_{\rm slope}^{2})^{\perp}$. We also endow the (even-degree) Mukai $F$-isocrystal
@@ 5.2. De Rham cohomology
where the ranks of the filtered pieces increase as $0,1,21,22.$ Using the perfect duality, we have $Fil^{1}_{\rm Hdg}=(Fil_{\rm Hdg}^{2})^{\perp}$. We also obtain from the above that
@@ 5.3. Proofs of the main results for K3 surfaces
${\bigoplus_{i=0}^{2}H_{\rm dR}^{2i}(X_{\operatorname{can},\eta}/K)(i)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\bigoplus_{i=0}^{2}H_{\rm cris}^{2i}(X/K)(i)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\Phi_{E}^{\rm cris}}$${\bigoplus_{i=0}^{2}H_{\rm dR}^{2i}(Y_{\operatorname{can},\eta}/K)(i)}$${\bigoplus_{i=0}^{2}H_{\rm cris}^{2i}(Y/K)(i)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$
@@
Since $E$ defines a derived equivalence $\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(X)\simeq\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(Y)$, we have
@@
A weaker version of the above can be obtained as follows as well. Let $X,Y$ be derived equivalent K3 surfaces over $k$. We know from [LO15] that $Y\simeq M_{v}(X)$ can be written as a certain fine moduli space of stable sheaves on $X$. Then, the relative moduli space $\mathcal{M}_{v}(X_{\operatorname{can}})$ over $W(k)$ is fine and Fourier-Mukai equivalent to $X_{\operatorname{can}}$ via the universal stable sheaf (Corollary 2.7). Proposition 5.5 implies that $Y_{\operatorname{can}}\simeq\mathcal{M}_{v}(X_{\operatorname{can}})$, so $X_{\operatorname{can}},Y_{\operatorname{can}}$ are relative Fourier-Mukai partners. This argument, however, does not show that arbitrary Fourier-Mukai kernels can be lifted.
@@ 6. Sample application to a result in [HLT20]
    AA is a degree 2 étale cyclic cover of $E\times_{k}F$.
@@
    AA is a degree 3 étale cyclic cover of $E\times_{k}F$, and $F$ has an automorphism of order 3.
@@ Appendix A Crystalline and de Rham Chern classes
${K_{0}(\operatorname{D}_{\operatorname{perf}}(\mathcal{X}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c_{\rm total,\rm dR}}$${K_{0}(\operatorname{D}_{\operatorname{perf}}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c_{\rm total,\rm cris}}$${H^{*}_{\rm dR}(\mathcal{X}/W)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${H^{*}_{\rm cris}(X/W)}$